# How to be an orthodox quantum mechanic

### v2

@@ Abstract
This work sets out to answer a single question: what is the orthodox interpretation of quantum mechanics? However, we adopt a different approach to that normally used. Rather than surveying physicists, or poring over the precise details of the thoughts of Bohr and Heisenberg, we review a collection of 42 textbooks on quantum mechanics, encompassing the most popular and prominent works of this nature. We then gauge their response to 12 propositions to build up a picture of exactly what is believed by an orthodox quantum mechanic. We demonstrate that this orthodoxy has many aspects of Copenhagen-like viewpoints, but also shows some interesting emerging deviations. Moreover, it is more nuanced than some reductive characterisations that condense the orthodoxy down to the ontological primacy of the quantum state. The revealed orthodoxy has two consistent pillars: measurement inherently disturbs quantum states and these states refer to individual instances, not ensembles. More fully, it entails that individual particles exist in wave-like super-positions and present particle behaviours only when forced to by outside influences. The act of measuring such a system inherently induces random changes in its state, manifesting as a form of measurement error that corresponds to the uncertainty principle. This implies that measurement does not reveal underlying values of quantum properties.
@@ 1 Introduction
How do scientists understand quantum mechanics? The most popular interpretation of quantum mechanics (QM) is usually found in surveys to be “Copenhagen interpretation” [1, 2]. However, the definition of this viewpoint is notoriously vague [3, 4, 5, 6], its characterisation tending to descend into hyperfine readings of Bohr and Heisenberg [4].
@@
Max Jammer sets out the quantum formalism as a set of 5 postulates, all of which are necessary to an orthodox or Copenhagen-like view [5]. These include the completeness of the state and the projection postulate. To produce something like a Copenhagen interpretation, this is supplemented by a kind of positivist objection to ascribing as real anything that cannot be directly measured, generating what Shimony [7] described as “doing epistemology without ontology”. This all results in a similar view to Home, that it is the primacy of the quantum state that typifies the orthodox view. Interestingly, Stapp argues [4] that a Copenhagen viewpoint does not see the quantum state as describing the evolution of the actual system in question, only its probabilities. This seems somewhat difficult to square with the notion that the state is the most complete specification of the real situation, without some explanation of why additional information is inaccessible. Home [6] argues that the primary characteristic of an orthodox viewpoint is that the state is framed as a complete description of a quantum system. Thus, this encapsulates both Copenhagen-like and ensembles-based theories. This latter inclusion stems from a claim that ensemble theories are internally incomplete unless they assume completeness of the quantum state (otherwise they must supply some hidden variables to assign definite values to measurable properties). Home refers to Feyeraband’s encapsulation of the standard position as being that quantum systems possess no definite properties over and above those that can be derived from the state itself [8].
@@
These points reveal a tension in descriptions of a quantum orthodoxy: a Copenhagen-like view wants to claim a “completeness” or “finality” for QM (the “end of the road hypothesis” as Popper terms it [9]). However, it also contains a strong degree of instrumentalism, as acknowledged by both Home and Jammer. This latter aspect would deny any ontological aspects to the theory. It follows then that the state is merely a prescription for the prediction of experimental outcomes. It seems clear that a follower of an allegedly Copenhagen viewpoint could pick and choose between these two, somewhat opposed, strands of thought.
@@
Rather than characterise this Copenhagen model, it seems more productive to instead determine what the consensus within the physics community actually is. With a particular focus on answers to questions about quantum phenomena, along the lines of those used in [1]. Existing surveys of physicists tend to reveal a very fragmented community [1, 2], instead we will look for a quantum mechanical orthodoxy in the many treatises that have followed from the foundational works usually analysed to define a Copenhagen interpretation.
@@
The target of our survey will be textbooks. The reason for this is that these works are presented for the education of a new generation of scientists. Thus, they are discussions of what we generally believe to be the ‘established facts’ of a given field. The key question is would the ‘established facts’ include an interpretive viewpoint? Perhaps textbooks function merely to convey a mathematical framework to their reader, with any interpretive remarks being negligible asides. This suggestion might have some validity, as a student of science can very often develop their intuition and understanding of a subject via that of the mathematics representing it. However, this is not straight-forwardly so in the realm of quantum mechanics. Evidenced by the proliferation of wildly different interpretations sharing the same (or very similar) mathematics. Thus, it is insufficient to pass on only a common mathematical core in this subject area. After all, most physics students will sketch out strikingly similar explanations of various unusual quantum results, even if their later work is done in deeply unrelated fields. They will also not merely bore a causal interlocutor with a barrage of equations. Where then would a student of quantum mechanics be exposed to an orthodoxy if not their textbooks? Perhaps in popular science media? However, one suspects that such media would itself be based on textbooks, rather than wade through the mire of competing interpretations in the active literature. An aspect of this can be seen in the fact that more recent textbooks devote more space to the interpretation of quantum phenomena (many have this as a separate chapter even). In addition, while working physicists may later deviate from the interpretations presented in reference works, we must acknowledge that their trajectory will be influenced by the initial conditions. Hence, the dominance of Copenhagen-like presentations in textbooks might lead to unorthodox interpretations which share many characteristics of this ‘reference orthodoxy’. An example of this would be the insistence that a quantum state does not apply merely to ensembles (as will be seen this is one of the most common assertions in textbooks). This is inherited by the vast majority of modern unorthodox interpretations (e.g. Bohmian mechanics [10], quantum Bayesianism [11], consistent histories [12], and relational quantum mechanics [13]), despite it not being necessary to the formalism [14]. Thus, even the beliefs of established researchers in quantum interpretation may be strongly correlated with the orthodoxy found in textbooks. It is therefore suggested that one cannot dismiss the interpretive remarks made in reference works on this subject. They fulfil a vital function in passing along a common understanding of quantum mechanics that is insufficiently conveyed by the mathematical formalism alone.
@@
To determine an orthodoxy empirically, we require some kind of framework of positions that might be part of the orthodoxy. Once again, this injects subjectivity into our analysis, as it is not clear how we might choose these positions. We settle on a set of 12 propositions, divided into three categories. These are selected to at least cover the major assertions of a Copenhagen-like framework and show strong overlap with surveys [1]. The first category explores how we understand the quantum state, with the propositions: “the state is ontologically complete”, “the state does not merely describe an ensemble of identically prepared systems”, and “the state collapses upon measurement”. The second category revolves around the nature of particles with the propositions: “individual particles have wave-like properties”, “particle-like entities are described by a wave-particle duality”, “particles cannot simultaneously possess definite values of non-commuting observables”, “particles do not follow classical-like trajectories”, and “there is no adequate single picture of particle or wave”. Finally, we turn to the question of why quantum mechanics predicts statistics only. Here we have the propositions: “measurement inherently disturbs quantum systems”, “measurement creates the observed definite values”, “the uncertainty principle results from measurement disturbance”, and “the limitation to prediction of probabilities results from the uncertainty principle”. These will all be elaborated on in later sections of this paper. Each work is scrutinised, and a trinary value is assigned to their agreement with each proposition (an in-between value is used when agreement can be strongly inferred for consistency with other positions taken). It can then be determined to what degree the works assent to each proposition on average and, by analysing the statistics, a set of propositions that represent a consensus viewpoint can be extracted (should one exist).
@@ 2 Propositions of an orthodoxy
Defining all aspects of an interpretative orthodoxy is a difficult proposition. The approach taken here is to define three sub-categories of hopefully fundamental aspects of the interpretation of the quantum formalism, these being the nature of the quantum state, that of particles, and why our predictions are all statistical. Within each category we include propositions that source works can either agree with or dissent from. The selection of these questions was designed partly apriori and partly by assertions encountered in some works. We reiterate that the choice of textbooks over research articles is made because we feel that the most distilled consensus is to be found in textbooks, which typically are formulated around what is considered “established knowledge” within a given field. It is entirely possible that this question set is incomplete (or contains some redundancy).
We also note that our choices cover questions asked of working physicists in surveys [1, 2]. Notably, the aforementioned [1] also includes “are values are defined prior to measurement?” (“measurement creates values”), “is randomness inherent in QM?” (strong overlap with “ontological completeness”), “where is the electron in a hydrogen orbital?” (“no trajectories”), “is macroscopic superposition possible?” (linked to “the state doesn’t just describe an ensemble”), and a question on the solution of the measurement problem is covered by the entire set of questions on why QM is statistical. However, we do not include questions about the role of the observer or the implications of Bell inequality violations. These latter points are highly contentious in the literature and do not receive consistent coverage in textbooks. The cases of incomplete overlap are down to the luxury of expositive works like textbooks. In that, our questions can be simpler, more numerous, and less nuanced than those in [1, 2], as they need not be processed in real time by the respondent.
@@ 2.2.1 Particles are wave-like
Individual particles exhibit wave behaviours such as diffraction and interference. The waves do not merely appear in the statistics of quantum systems. Viewpoints that might deny this proposition would be a de Broglie-Bohm interpretation [10] (where wave and particle co-exist as linked entities) or an ensemble one (where the waves are merely statistical patterns).
@@ 2.2.5 The uncertainty principle applies to single particles
A common joke about QM goes: Heisenberg is stopped by traffic police who ask him “Do you know how fast you were going?”. His reply is “No, but I know exactly where I am”. This is the essence of this proposition. That uncertainty applies to individual realisations of quantum systems. That is, if I measure the position of a particle I cannot also know its momentum simultaneously. More generally: systems cannot possess definite values of non-commuting observables simultaneously. This is opposed to a statistical understanding which recognises that the uncertainty principle is derived via standard deviations over an ensemble of measurements performed on similarly prepared systems [9].
2.3 Why are our predictions statistical?
@@ 2.3.1 Measurement inherently disturbs the system
Measurement inherently changes the state of a system while measuring it (except in the case of eigenstates). A viewpoint that denies this would note that non-interactive measurements exist in QM or that measurement has no abstract definition.
@@ 2.3.3 Uncertainty leads to probabilities
This proposition regards the uncertainty principle as a fundamental feature of nature. Particularly, this also requires single-particle uncertainty. The consequences of this is that measurements have probabilistic results, even though quantum mechanics is not formulated as fundamentally statistical. One might deny this if one recognises that the uncertainty principle is a consequence of quantum rules, rather than a postulate, or if one notes the uncertainty principle is formulated in terms of statistical observables only.
2.3.4 Measurement creates observed values
The observed values of quantum properties are the result of an interaction between apparatus and system that alters the measured property. In this sense, they do not reflect underlying pre-measurement values in cases that are not eigenstates of the observable being measured. This stands in opposition to the usual scientific idea of measurement: that we are generally discovering the values of pre-existing properties. This assertion can take two forms: a strong form, where it is asserted that quantum systems do not actually possess definite properties till measured, and a weaker form where measurement merely changes pre-existing values. Both of these result in the pre-measurement value being indeterminate and potentially irrelevant.
@@ 3 Sample of studied works
To perform this study we examine the $42$ popular textbooks listed in Table 1. These range in date of first publication from 1930 to 2025. This large time span is of utility, as it allows us to determine whether there is any evolution within the consensus view of quantum mechanics. Additionally, we consider subdivision of these works into graduate ($N=17$) and undergraduate ($N=25$), as determined from the statements of the target audience in the work itself. These works were chosen to encompass the Wikipedia list on ‘Notable quantum mechanics textbooks’ [^2] https://en.wikipedia.org/wiki/List_of_textbooks_on_classical_mechanics_and_quantum_mechanics ($N=18$) as well as to include as large a sample as feasible. To detect variation over time we also consider three subsamples: works pre-1980 ($N=15$), post-1980 ($N=27$), and post-2000 ($N=17$). Note that Omnés work is included, even though it is a textbook that advocates against the orthodoxy in favour of “consistent histories”, as it provides its own description of the orthodox positions which we use for our study. In the case of [15], this thoughtfully discusses some foundational issues without expressing a direct opinion, here we substitute what the author characterises as the “conventional” approach. Note that page references for [16] are in reference to the online lecture notes on which it is based [^3] https://www.damtp.cam.ac.uk/user/tong/books/quantum.html, due to the recency of the text. We will also make reference to a set of 4 works whom we do not include in any of the samples: [6, 17, 18, 14]. The first is excluded as it is a discussion of interpretations, but will provide a useful point of comparison. The second two are analysed out of curiosity, but are not targeted squarely as part of a physics syllabus. The final one will be analysed and found to be entirely outside the orthodoxy on almost every point of contention (so it is treated as an outlier).
@@
Lecture notes on quantum mechanics D. Tong 2025 [16] U N
Basic quantum mechanics K. Tamvakis 2019 [19] U N
A first introduction to quantum physics P. Kok 2018 [20] U N
Lectures on quantum mechanics S. Weinberg 2013 [21] G N
The physics of quantum mechanics J. Binney & D. Skinner 2013 [22] U Y
Quantum concepts in physics M. Longair 2013 [23] G N
Quantum mechanics D. McIntyre et al 2012 [24] U Y
Quantum mechanics: theory and experiment M. Beck 2012 [25] U N
Quantum mechanics: a new introduction K. Konishi & G. Paffuti 2009 [26] U N
Quantum mechanics for scientists and engineers D. Miller 2008 [27] U N
Quantum mechanics: a textbook for undergraduates M. Jain 2007 [28] U N
Introduction to quantum mechanics H. Müller-Kirsten 2006 [29] U Y
Quantum physics M. Le Bellac 2006 [30] G N
Lectures on quantum mechanics J. Basdevant 2005 [31] G N
Introduction to quantum mechanics A. Philips 2003 [32] U N
Quantum Mechanics: Concepts and Applications N. Zettili 2001 [33] U Y
Elements of quantum mechanics M. Fayer 2001 [34] U N
Understanding quantum mechanics R. Omnés 1999 [35] U N
Particles behave like waves T. Moore 1998 [36] U N
Introduction to quantum mechanics D. Griffiths 1995 [37] U Y
Quantum theory A. Peres 1993 [38] G Y
Quantum mechanics F. Mandl 1992 [39] U N
A modern approach to quantum mechanics J. Townsend 1992 [40] U Y
Quantum mechanics: an introduction W. Greiner 1989 [41] U N
Quantum mechanics A. Rae 1986 [15] U N
Modern quantum mechanics J. Sakurai 1985 [42] G Y
Principles of Quantum mechanics R. Shankar 1980 [43] G Y
An introduction to quantum physics A. French & E. Taylor 1978 [44] U Y
Quantum mechanics C. Cohen-Tannoudji 1977 [45] G Y
Quantum physics of atoms, molecules, solids, nuclei and particles R. Eisberg & R. Resnick 1974 [46] U Y
Quantum physics S. Gasiorowicz 1974 [47] U Y
Feynman lectures on physics volume 3 R. Feynman 1965 [48] U Y
Quantum mechanics A. Davydov 1965 [49] G Y
Elementary quantum mechanics P. Fong 1962 [50] U N
Quantum mechanics E. Merzbacher 1961 [51] G N
Introduction to quantum mechanics R. Dicke & J. Wittke 1960 [52] G N
Quantum mechanics A. Messiah 1959 [53] G N
Quantum mechanics: non-relativistic theory L. Landau & L. Lifshitz 1958 [54] G Y
Mathematical foundations of quantum mechanics J. von Neumann 1955 [55] G Y
Quantum theory D. Bohm 1951 [56] G N
Quantum mechanics L. Schiff 1949 [57] G N
The principles of quantum mechanics P. Dirac 1930 [58] G Y
Quantum mechanics: a modern development L. Ballentine 1998 [14] G N
Quantum computation and quantum information M. Nielsen & I. Chuang 2000 [18] G N
An introduction to quantum theory K. Hannabuss 1997 [17] G N
Conceptual foundations of quantum physics D. Home 1997 [6] - N

### v3

@@ Abstract
This work sets out to answer a single question: what is the orthodox interpretation of quantum mechanics? However, we adopt a different approach to that normally used. Rather than surveying physicists, or poring over the precise details of the thoughts of Bohr and Heisenberg, we review a collection of 42 textbooks on quantum mechanics, encompassing the most popular and prominent works of this nature. We then gauge their response to 13 propositions to build up a picture of exactly what is believed by an orthodox quantum mechanic. We demonstrate that this orthodoxy has many aspects of Copenhagen-like viewpoints, but also shows some interesting emerging deviations. Moreover, it is more nuanced than some reductive characterisations that condense the orthodoxy down to the ontological primacy of the quantum state. The revealed orthodoxy has three consistent pillars: measurement inherently disturbs quantum states, these states refer to individual instances, not ensembles, and quantum systems do not have definite properties prior to measurement. More fully, it entails that individual particles exist in wave-like super-positions and present particle behaviours only when forced to by outside influences. The act of measuring such a system inherently induces random changes in its state, manifesting as a form of measurement error that corresponds to the uncertainty principle. This implies that measurement does not reveal underlying values of quantum properties.
@@ 1 Introduction
How do scientists understand quantum mechanics? The most popular interpretation of quantum mechanics (QM) is usually found in surveys to be “Copenhagen interpretation” [sivasundaram2016surveyingattitudesphysicistsconcerning, survey2025]. However, the definition of this viewpoint is notoriously vague [hanson1959, stapp1972, Jammer1974-JAMTPO-10, Home1997], its characterisation tending to descend into hyperfine readings of Bohr and Heisenberg [stapp1972].
@@
Max Jammer sets out the quantum formalism as a set of 5 postulates, all of which are necessary to an orthodox or Copenhagen-like view [Jammer1974-JAMTPO-10]. These include the completeness of the state and the projection postulate. To produce something like a Copenhagen interpretation, this is supplemented by a kind of positivist objection to ascribing as real anything that cannot be directly measured, generating what Shimony [shimony1993] described as “doing epistemology without ontology”. This all results in a similar view to Home, that it is the primacy of the quantum state that typifies the orthodox view. Interestingly, Stapp argues [stapp1972] that a Copenhagen viewpoint does not see the quantum state as describing the evolution of the actual system in question, only its probabilities. This seems somewhat difficult to square with the notion that the state is the most complete specification of the real situation, without some explanation of why additional information is inaccessible.
Home [Home1997] argues that the primary characteristic of an orthodox viewpoint is that the state is framed as a complete description of a quantum system. Thus, this encapsulates both Copenhagen-like and ensembles-based theories. This latter inclusion stems from a claim that ensemble theories are internally incomplete unless they assume completeness of the quantum state (otherwise they must supply some hidden variables to assign definite values to measurable properties). Home refers to Feyeraband’s encapsulation of the standard position as being that quantum systems possess no definite properties over and above those that can be derived from the state itself [feyeraband1962].
@@
These points reveal a tension in descriptions of a quantum orthodoxy: a Copenhagen-like view wants to claim a “completeness” or “finality” for QM (the “end of the road hypothesis” as Popper termed it [Popper1982-POPQTA]). However, it also contains a strong degree of instrumentalism, as acknowledged by both Home and Jammer. This latter aspect would deny any ontological aspects to the theory. It follows then that the state is merely a prescription for the prediction of experimental outcomes. It seems clear that a follower of an allegedly Copenhagen viewpoint could pick and choose between these two, somewhat opposed, strands of thought.
@@
Rather than characterise this Copenhagen model, it seems more productive to instead determine what the consensus within the physics community actually is. With a particular focus on answers to questions about quantum phenomena, along the lines of those used in [sivasundaram2016surveyingattitudesphysicistsconcerning]. Existing surveys of physicists tend to reveal a very fragmented community [sivasundaram2016surveyingattitudesphysicistsconcerning, survey2025], instead we will look for a quantum mechanical orthodoxy in the many treatises that have followed from the foundational works usually analysed to define a Copenhagen interpretation.
@@
The target of our survey will be textbooks. The reason for this is that these works are presented for the education of a new generation of scientists. Thus, they are discussions of what we generally believe to be the ‘established facts’ of a given field. The key question is would the ‘established facts’ include an interpretive viewpoint? Perhaps textbooks function merely to convey a mathematical framework to their reader, with any interpretive remarks being negligible asides. This suggestion might have some validity, as a student of science can very often develop their intuition and understanding of a subject via that of the mathematics representing it. However, this is not straight-forwardly so in the realm of quantum mechanics. Evidenced by the proliferation of wildly different interpretations sharing the same (or very similar) mathematics. Thus, it is insufficient to pass on only a common mathematical core in this subject area. After all, most physics students will sketch out strikingly similar explanations of various unusual quantum results, even if their later work is done in deeply unrelated fields. They will also not merely bore a causal interlocutor with a barrage of equations. Where then would a student of quantum mechanics be exposed to an orthodoxy if not their textbooks? Perhaps in popular science media? However, one suspects that such media would itself be based on textbooks, rather than wade through the mire of competing interpretations in the active literature. An aspect of this can be seen in the fact that more recent textbooks devote more space to the interpretation of quantum phenomena (some have this as a separate chapter even). In addition, while working physicists may later deviate from the interpretations presented in reference works, we must acknowledge that their trajectory will be influenced by the initial conditions. Hence, the dominance of Copenhagen-like presentations in textbooks might lead to unorthodox interpretations which share many characteristics of this ‘reference orthodoxy’. An example of this would be the insistence that a quantum state does not apply merely to ensembles (as will be seen this is one of the most common assertions in textbooks). This is inherited by the vast majority of modern unorthodox interpretations (e.g. Bohmian mechanics [bohm1982], quantum Bayesianism [Fuchs_2013], consistent histories [sep-qm-consistent-histories], and relational quantum mechanics [Rovelli_1996]), despite it not being necessary to the formalism [ballentine1998quantum]. Thus, even the beliefs of established researchers in quantum interpretation may be strongly correlated with the orthodoxy found in textbooks. It is therefore suggested that one cannot dismiss the interpretive remarks made in reference works on this subject. They fulfil a vital function in passing along a common understanding of quantum mechanics that is insufficiently conveyed by the mathematical formalism alone.
@@
To determine an orthodoxy empirically, we require some kind of framework of positions that might be part of the orthodoxy. Once again, this injects subjectivity into our analysis, as it is not clear how we might choose these positions. We settle on a set of 13 propositions, divided into three categories. These are selected to at least cover the major assertions of a Copenhagen-like framework and show strong overlap with surveys [sivasundaram2016surveyingattitudesphysicistsconcerning]. The first category explores how we understand the quantum state, with the propositions: “the state is ontologically complete”, “the state does not merely describe an ensemble of identically prepared systems”, and “the state collapses upon measurement”. The second category revolves around the nature of particles with the propositions: “individual particles have wave-like properties”, “fundamental wave-like behaviour is reflected by the wave-particle duality”, “particles cannot simultaneously possess definite values of non-commuting observables”, “particles do not follow classical-like trajectories”, and “there is no adequate single picture of particle or wave”. Next, we turn to the question of why quantum mechanics predicts statistics only. Here we have the propositions: “measurement inherently disturbs quantum systems”, “quantum systems have no property values till measured”, “the uncertainty principle results from measurement disturbance”, and “the limitation to prediction of probabilities results from the uncertainty principle”. Finally, we have “other interpretations of quantum mechanics are irrelevant”, which doesn’t fall into any of the categories. These will all be elaborated on in later sections of this paper. Each work is scrutinised, and a trinary value is assigned to their agreement with each proposition (an in-between value is used when agreement can be strongly inferred for consistency with other positions taken). It can then be determined to what degree the works assent to each proposition on average and, by analysing the statistics, a set of propositions that represent a consensus viewpoint can be extracted (should one exist).
@@ 2 Propositions of an orthodoxy
Defining all aspects of an interpretative orthodoxy is a difficult proposition. The approach taken here is to define three sub-categories of hopefully fundamental aspects of the interpretation of the quantum formalism, these being the nature of the quantum state, that of particles, and why our predictions are probabilistic. Within each category we include propositions that source works can either agree with or dissent from. The selection of these questions was designed partly apriori and partly by assertions encountered in some works. We reiterate that the choice of textbooks over research articles is made because we feel that the most distilled consensus is to be found in textbooks, which typically are formulated around what is considered “established knowledge” within a given field. It is entirely possible that this question set is incomplete (or contains some redundancy).
We also note that our choices cover questions asked of working physicists in surveys [sivasundaram2016surveyingattitudesphysicistsconcerning, survey2025]. Notably, the aforementioned survey [sivasundaram2016surveyingattitudesphysicistsconcerning] also includes “are values are defined prior to measurement?” (“no pre-measurement values”), “is randomness inherent in QM?” (strong overlap with “ontological completeness”), “where is the electron in a hydrogen orbital?” (“no trajectories”), “is macroscopic superposition possible?” (linked to “the state doesn’t just describe an ensemble”), and a question on the solution of the measurement problem is covered by the entire set of questions on why QM is probabilistic. However, we do not include questions about the role of the observer or the implications of Bell inequality violations. These latter points are highly contentious in the literature and do not receive consistent coverage in textbooks. The cases of incomplete overlap are down to the luxury of expositive works like textbooks. In that, our questions can be simpler, more numerous, and less nuanced than those in [sivasundaram2016surveyingattitudesphysicistsconcerning, survey2025], as they need not be processed in real time by the respondent.
@@ 2.2.1 Particles are wave-like
Individual particles exhibit wave behaviours such as diffraction and interference. The waves do not merely appear in the statistics of quantum systems. Viewpoints that might deny this proposition would be a de Broglie-Bohm interpretation [bohm1982] (where wave and particle co-exist as linked entities) or an ensemble one (where the waves are merely statistical patterns).
@@ 2.2.5 The uncertainty principle applies to single particles
A common joke about QM goes: Heisenberg is stopped by traffic police who ask him “Do you know how fast you were going?”. His reply is “No, but I know exactly where I am”. This is the essence of this proposition. That uncertainty applies to individual realisations of quantum systems. That is, if I measure the position of a particle I cannot also know its momentum simultaneously. More generally: systems cannot possess definite values of non-commuting observables simultaneously. This is opposed to a statistical understanding which recognises that the uncertainty principle is derived via standard deviations over an ensemble of measurements performed on similarly prepared systems [Popper1982-POPQTA].
2.3 Why are our predictions probabilistic?
@@ 2.3.1 Measurement inherently disturbs the system
Measurement inherently changes the state of a system while measuring it (except in the case of eigenstates). A viewpoint that denies this would note that non-interactive measurements exist in QM, or that measurement has no abstract definition.
@@ 2.3.3 Uncertainty leads to probabilities
This proposition regards the uncertainty principle as a fundamental feature of nature. Particularly, this also requires single-particle uncertainty. The consequences of this is that measurements have probabilistic results, even though quantum mechanics is not formulated as fundamentally probabilistic. One might deny this if one recognises that the uncertainty principle is a consequence of quantum rules, rather than a postulate, or if one notes the uncertainty principle is formulated in terms of statistical observables only.
2.3.4 There are no pre-measurement values
Except for those observable eigenstates, quantum systems do not have definite properties that we uncover via measurement. Instead, the value comes to exist at the point of measurement. A primary signature of this is how the Born rule is described. If a text states that $|\psi(x)|^{2}$ is the probability of finding the system with property value $x$, then it affirms this proposition. However, this alone is not enough, we only infer assent if there is other supportive text as well. A dissenting text would instead state the Born rule as giving the probability of a system having property value $x$. This appears to be a rather fine point of semantics. However, almost all the surveyed texts use precisely the same wording: finding, not having. The ones that don’t use “finding”, seem to do so deliberately, and it correlates with more direct assertions that pre-measurement values exist.
2.4 Other propositions
These don’t fit into the above categories.
2.4.1 There is only the orthodoxy
Only the orthodox presentation of quantum mechanics is worthwhile or serious. An affirming text either mentions no other interpretations or dismisses them all as unconvincing or disproven. A textbook that notes it is presenting the orthodox/conventional/standard approach but admits this isn’t definitive will be taken as dissenting.
@@ 3 Sample of studied works
To perform this study we examine the $42$ popular textbooks listed in Table 1. These range in date of first publication from 1930 to 2025. This large time span is of utility, as it allows us to determine whether there is any evolution within the consensus view of quantum mechanics. Additionally, we consider subdivision of these works into graduate ($N=17$) and undergraduate ($N=25$), as determined from the statements of the target audience in the work itself. These works were chosen to encompass the Wikipedia list on ‘Notable quantum mechanics textbooks’ [^2] https://en.wikipedia.org/wiki/List_of_textbooks_on_classical_mechanics_and_quantum_mechanics ($N=18$) as well as to include as large a sample as feasible. To detect variation over time we also consider three subsamples: works pre-1980 ($N=15$), post-1980 ($N=27$), and post-2000 ($N=17$). Note that Omnés work is included, even though it is a textbook that advocates against the orthodoxy in favour of “consistent histories”, as it provides its own description of the orthodox positions which we use for our study. In the case of [rae2002], this thoughtfully discusses some foundational issues without expressing a direct opinion, here we substitute what the author characterises as the “conventional” approach. We will also make reference to a set of 4 works whom we do not include in any of the samples: [Home1997, Hannabuss:1999fa, nielsen2010, ballentine1998quantum]. The first is excluded as it is a discussion of interpretations, but will provide a useful point of comparison. The second two are analysed out of curiosity, but are not targeted squarely as part of a physics syllabus. The final one will be analysed and found to be entirely outside the orthodoxy on almost every point of contention (so it is treated as an outlier).
@@
Lecture notes on quantum mechanics D. Tong 2025 [tong2025] U N
Basic quantum mechanics K. Tamvakis 2019 [tamvakis2019] U N
A first introduction to quantum physics P. Kok 2018 [kok2023] U N
Lectures on quantum mechanics S. Weinberg 2013 [weinberg2013] G N
The physics of quantum mechanics J. Binney & D. Skinner 2013 [binney2013physics] U Y
Quantum concepts in physics M. Longair 2013 [longair_2013] G N
Quantum mechanics D. McIntyre et al 2012 [mcintyre2012quantum] U Y
Quantum mechanics: theory and experiment M. Beck 2012 [beck2012] U N
Quantum mechanics: a new introduction K. Konishi & G. Paffuti 2009 [Konishi:2009qva] U N
Quantum mechanics for scientists and engineers D. Miller 2008 [miller2010] U N
Quantum mechanics: a textbook for undergraduates M. Jain 2007 [jain2017] U N
Introduction to quantum mechanics H. Müller-Kirsten 2006 [muller-kisten2006] U Y
Quantum physics M. Le Bellac 2006 [bellac2011quantum] G N
Lectures on quantum mechanics J. Basdevant 2005 [basdevant2023] G N
Introduction to quantum mechanics A. Philips 2003 [phillips2013introduction] U N
Quantum Mechanics: Concepts and Applications N. Zettili 2001 [zettili2001] U Y
Elements of quantum mechanics M. Fayer 2001 [fayer2001] U N
Understanding quantum mechanics R. Omnés 1999 [Omnes1999-OMNUQM] U N
Particles behave like waves T. Moore 1998 [moore1998] U N
Introduction to quantum mechanics D. Griffiths 1995 [Griffiths2004Introduction] U Y
Quantum theory A. Peres 1993 [peres1993] G Y
Quantum mechanics F. Mandl 1992 [mandl2013quantum] U N
A modern approach to quantum mechanics J. Townsend 1992 [townsend2012modern] U Y
Quantum mechanics: an introduction W. Greiner 1989 [greiner1989] U N
Quantum mechanics A. Rae 1986 [rae2002] U N
Modern quantum mechanics J. Sakurai 1985 [Sakurai:1167961] G Y
Principles of Quantum mechanics R. Shankar 1980 [Shankar:102017] G Y
An introduction to quantum physics A. French & E. Taylor 1978 [french1978] U Y
Quantum mechanics C. Cohen-Tannoudji 1977 [Cohen-Tannoudji:101367] G Y
Quantum physics of atoms, molecules, solids, nuclei R. Eisberg & R. Resnick 1974 [eisberg1974] U Y
and particles
Quantum physics S. Gasiorowicz 1974 [gasiorowicz2007quantum] U Y
Feynman lectures on physics volume 3 R. Feynman 1965 [feynman1977feynman] U Y
Quantum mechanics A. Davydov 1965 [davydov1965] G Y
Elementary quantum mechanics P. Fong 1962 [fong2013elementary] U N
Quantum mechanics E. Merzbacher 1961 [merzbacher1998] G N
Introduction to quantum mechanics R. Dicke & J. Wittke 1960 [dicke1960] G N
Quantum mechanics A. Messiah 1959 [messiah1999quantum] G N
Quantum mechanics: non-relativistic theory L. Landau & L. Lifshitz 1958 [Landau1981Quantum] G Y
Mathematical foundations of quantum mechanics J. von Neumann 1955 [vonneumann1955] G Y
Quantum theory D. Bohm 1951 [bohm1951] G N
Quantum mechanics L. Schiff 1949 [schiff1955] G N
The principles of quantum mechanics P. Dirac 1930 [Dirac1930-DIRTPO] G Y
Quantum mechanics: a modern development L. Ballentine 1998 [ballentine1998quantum] G N
Quantum computation and quantum information M. Nielsen & I. Chuang 2000 [nielsen2010] G N
An introduction to quantum theory K. Hannabuss 1997 [Hannabuss:1999fa] G N
Conceptual foundations of quantum physics D. Home 1997 [Home1997] - N